OpenStudy (anonymous):

What is the simplified form of 1/x^3/3/x^2

OpenStudy (mertsj):

\[\frac{1}{x^3}\times\frac{x^2}{3}=\frac{1}{3x}\]

OpenStudy (anonymous):

is that all you do

OpenStudy (mertsj):

That's as simple as it gets.